a stone is thrown vertically upwards with an initial speed of u, its vertically displacement s after an elapsed time t is given by the formular s= ut - gt^2/2 where g is the acceleration due to gravity air resistance is ignored, calculate the value of s over a period of 12.3 seconds at intervals of 0.1 seconds
. and plot the distance time graph over this period by writing matlab/ octave code that implements and solves rhis problem.
Answer #1

Uniforn_motion.m+ t 0:0.1:12.3; 1 2 g-9.8; u=50m/s u=70; %say, 3 4 plot (t, s); 5 grid on 6 title(distance vs time plot xladistance vs time plot 250 200 150 100 50 12 10 14 time () in seconds

you can choose any value of u. the graph will change accordingly. in the code section the line in green should have been %say u=70 m/s instead of 50m/s, since i plotted with u=70. ignore that as a typing mistake